1.108.1 Use and manage local system documentation Weight 4

Linux Professional Institute Certification — 102

Geoffrey Robertson ge@ffrey.com

               Nick Urbanik nicku@nicku.org

2005 July

Documentation

2.108.1

Use and Manage Local System Documentation [4]

2.108.2 Find Linux Documentation on the Internet [3]

2.108.5 Notify Users on System Related Issues [1]

Use and Manage Local System Documentation

Candidates should be able to use and administer the man facility and the material in /usr/share/doc/.

This objective includes finding relevant man pages, searching man page sections, finding commands and man pages related to them, and configuring access to man sources and the man system.

It also includes using system documentation stored in /usr/share/doc/ and determining what documentation to keep in /usr/share/doc/.

Use and Manage Local System Documentation

MANPATH
man
apropos
whatis

\underline{2}

Use and Manage Local System Documentation

Using whatis

1 whatis searches the manual page names and displays the manual page descrip tions of any name matched.

2

$ whatis nmap ↩
nmap   (1)  - Network exploration tool
                and security scanner

3 (in Debian Woody but not RH9.0)

-r, –regex     Interpret each name as
                 a regular expression.
$ whatis -r ^ifc ↩
ifconfig (8)   - configure a network interface
Using apropos

Each manual page has a short description available within it. apropos searches the descriptions for instances of keyword.

1

$ apropos nmap ↩
MapWindow (3tk)       - map or unmap a window
mmap (2)               - map or unmap files or d
munmap (2)             - map or unmap files or devi
nmap (1)               - Network exploration tool and
k_MapWindow (3tk)   - map or unmap a window
...

Using man

$ man man ↩
man(1)                                                                           man(1)
NAME
        man - format and display the on-line manual pages
        manpath - determine user search path for man pages
SYNOPSIS
        man  [-acdfFhkKtwW]  [–path]  [-m system] [-p string] [-C config_file]

\underline{3}

      [-M pathlist] [-P pager] [-S section_list] [section] name ...
DESCRIPTION
      man formats and displays the on-line manual pages.

Using man -k Equivalent to apropos -f Equivalent to whatis -K Search for the specified string in *all* man pages. Warning: this is probably very slow!

It helps to specify a section. (Just to give a rough idea, on my machine this takes about a minute per 500 man pages.)

The Manual Section Numbers 1. Executable programs or shell commands

2. System calls (functions provided by the kernel)

3. Library calls (functions within system libraries)

4. Special files (usually found in /dev)

5. File formats and conventions eg /etc/passwd

6. Games

7. Miscellaneous (including macro packages and conven tions), e.g. man(7), groff(7)

8. System administration commands (usually only for root)

9. Kernel routines [Non standard]

GNU info
File: info,  Node: Top,  Next: Getting Started,  Up: (dir)
Info: An Introduction
*********************
Info is a program, which you are using now, for reading
documentation of computer programs.  The GNU Project distributes most
of its on-line manuals in the Info format, so you need a program called
"Info reader" to read the manuals.  One of such programs you are using
now.

\underline{4}

   If you are new to Info and want to learn how to use it, type the
command `h' now.  It brings you to a programmed instruction sequence.
   To learn advanced Info commands, type `n' twice.  This brings you to
`Info for Experts', skipping over the `Getting Started' chapter.
* Menu:
* Getting Started::               Getting started using an Info reader.
* Advanced Info::                 Advanced commands within Info.
* Creating an Info File          How to make your own Info file.
–zz-Info: (info.gz)Top, 24 lines –Top—-*** Tags out of Date ***———–
Welcome to Info version 4.2. Type C-h for help, m for menu item.

GNU info Navigation Tab Move among hyperlinks Enter Follow hyperlinks d Return to the top ? List these info commands n Move to the "next" node of this node. p Move to the "previous" node of this node. u Move "up" from this node. q Quit Info altogether. h Show a primer for first time users. /string Search for string.

/usr/share/doc/

General information about packages and programs.

$ ls /usr/share/doc/ ↩
4Suite-0.11.1                      libtabe-0.2.6
a2ps-4.13b                         libtiff-3.5.7
abiword-1.0.4                      libtiff-devel-3.5.7
ac-archive-0.5.39                 libtool-1.4.3
acl                                 libungif-4.1.0
adjtimex-1.13                      libungif-devel-4.1.0
alchemist-1.0.26                  libunicode-0.4

/usr/share/doc/
$ ls /usr/share/doc/acl ↩
CHANGES.gz  LICENSE  PORTING  README
/usr/share/doc/
$ head /usr/share/doc/acl/README ↩
Access control list package README
__________________________________
See the file doc/INSTALL for build, installation and post-
install configuration steps.
Refer to the acl(5) manual page for general access control list
information and references to other related manual pages.
The extended attributes and access control list project for the
ext2 and ext3 filesystems is located at:
         http://acl.bestbits.at/
For more information and details on how to contribute to the
XFS project see the web pages at: http://oss.sgi.com/projects/xfs/
/usr/share/doc/
   $ mozilla /usr/share/doc/xfig/html/index.html ↩

The End

License Of This Document